OpenStudy (anonymous):

could someone show me the formula at a interest rate of 12% compounded quarterly looking to get the (apy)

jimthompson5910 (jim_thompson5910):

Use the formula APY = (1+r/n)^n - 1 where APY = annual percentage yield r = interest rate n = compounding frequency In this case APY = unknown r = 0.12 n = 4

OpenStudy (anonymous):

thank you

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

came up with answer 624 but is not giving me the apy

OpenStudy (anonymous):

did i miss something?

jimthompson5910 (jim_thompson5910):

APY = (1+r/n)^n - 1 APY = (1+0.12/4)^4 - 1 APY = (1+0.03)^4 - 1 APY = (1.03)^4 - 1 APY = 1.12550881 - 1 APY = 0.12550881

OpenStudy (anonymous):

lol

jimthompson5910 (jim_thompson5910):

So if the interest rate is 12%, compounded quarterly, then the APY is roughly 12.550881%

jimthompson5910 (jim_thompson5910):

so it's close to the interest rate (but slightly higher)

OpenStudy (anonymous):

thats even more complicated but nicely explained like that

OpenStudy (anonymous):

i like that

jimthompson5910 (jim_thompson5910):

you can read more about APY (aka effective interest rate) here http://en.wikipedia.org/wiki/Effective_interest_rate

OpenStudy (anonymous):

i was trying to get to that apy and thx for link now trying it thx jim

jimthompson5910 (jim_thompson5910):

yw

OpenStudy (anonymous):

now i can try my new scientific calculator lol using the above

jimthompson5910 (jim_thompson5910):

hopefully you'll get the same

OpenStudy (anonymous):

hope so

OpenStudy (anonymous):

yes i did thx

jimthompson5910 (jim_thompson5910):

ok great, yw

OpenStudy (anonymous):

you made that look easy, now that i have studied your variations i completly understand not bad for a 50 + student lol